OpenStudy (anonymous):

Solve x² - 12x + 1 = 0 by completing the square.

OpenStudy (zzr0ck3r):

For \(x^2+bx+c=0\) we have \((x+\frac{b}{2})^2=(\frac{b}{2})^2-c\)